OpenStudy (anonymous):

Evaluate the trig expression cos120 degrees

OpenStudy (anonymous):

\[\cos \left( 180-\theta \right)=-\cos \theta \]

jimthompson5910 (jim_thompson5910):

hint: 120 degrees is in quadrant 2 (Q2) the reference angle formula for Q2 is R = 180-x where x is the given angle and R is the reference angle

OpenStudy (anonymous):

so 180-120?

jimthompson5910 (jim_thompson5910):

yeah so the reference angle is 60 degrees

OpenStudy (anonymous):

Is that it?

jimthompson5910 (jim_thompson5910):

now evaluate cos(60)

OpenStudy (anonymous):

-0.95

jimthompson5910 (jim_thompson5910):

cos(120) = -cos(60) keep in mind cosine is negative in Q2

jimthompson5910 (jim_thompson5910):

make sure you are in degree mode

OpenStudy (anonymous):

-0.0166?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Sorry if it's coming out weird. I don't have a scientific calculator (it broke) so I'm using the google calculator

jimthompson5910 (jim_thompson5910):

type in "cos(60 degrees)" without quotes

OpenStudy (anonymous):

0.5!

jimthompson5910 (jim_thompson5910):

google defaults in radian mode so you have to explicitly tell it to be in degree mode

OpenStudy (anonymous):

\[\cos 60=\frac{ 1 }{ 2 }\]

jimthompson5910 (jim_thompson5910):

so your final answer should be -1/2

OpenStudy (anonymous):

Ok :) Thanks!

jimthompson5910 (jim_thompson5910):

yw